# Enabling Fast and Accurate Neutral Atom Readout through Image Denoising

### v1

@@ 3.2. Experimental setup and dataset
In this work, we explore the advantages of employing machine learning to improve the accuracy and speed of measuring qubits in a neutral atom quantum computer. The researchers use a grid of cesium atoms, each held in place by optical tweezers. To read out the state of each atom, they shine laser light that causes atoms in the “bright” state to fluorescence, while “dark” state atoms emit little to no light. The emitted light is captured by a sensitive camera through two optical paths: a bright, high-quality path used for generating accurate labels, and a dimmer, noisier path used for testing machine learning models.
The experimental setup is shown in Figure 3, uses an EMCCD camera to capture the photons emitted by the atom due to fluorescence. The primary path captures entire signal, i.e, photons emitted by the atom during the fluorescence, while the secondary path attenuates this signal by resulting in equivalent shorter exposure setting, leading to reduction in readout time. The difference is evident from the histograms shown in the Figure 3, where the long exposure images shows the histograms with better separation relative to short exposure images.
The authors developed two convolutional neural networks (CNNs) to classify the atom states from images. The first, CNN-site, looks at one atom at a time. The second, CNN-array, processes the full image and can account for light spillover (crosstalk) between neighboring atoms—an issue that worsens when atoms are spaced closely together.
They trained the CNNs on labeled data from the high-quality path and tested them on the noisy images. Compared to standard threshold-based methods, both CNNs achieved significantly lower error rates. CNN-site reduced errors by up to 56% in low-crosstalk settings, while CNN-array cut errors by up to 43% and crosstalk by 78% in tightly packed arrays relative to Gaussian thresholding.
Overall, the machine learning models allows faster readout while maintaining high accuracy—an important step toward real-time qubit measurement in scalable quantum systems.
@@ 3.3. Why are standard CNN models not enough?
CNNs are the de-facto models for image classification. In natural-image domains with rich semantic content and large training datasets, they perform remarkably well, even in the presence of moderate image noise, label noise, and variations in scale or rotation, as they learn robust, hierarchical representations of visual features.
CNNs operate by first detecting local, low-level features such as edges, corners, and textures in early convolutional layers. Deeper layers combine these primitives into higher-level features that capture complex shapes and semantic structure. This feature hierarchy allows CNNs to distinguish objects even under degraded imaging conditions, provided there is enough semantic variation in the data.
Neutral atom qubit readout images differ fundamentally from natural-image datasets. Each image consists of fluorescence from atoms arranged on a fixed lattice, where each site is either bright () or dark (). Short exposure measurements capture roughly ten times fewer photons than long-exposure images, resulting in significantly lower SNR, as shown in Figure 4(a)–(b). The corresponding pixel-intensity histograms in Figure 4(c)–(d) illustrates this effect of bright and dark sites overlapping heavily at short exposures, while long exposures yield well-separated distributions, making accurate state discrimination difficult for low-SNR images.
Unlike natural images, neutral atom readouts lack semantic structure and variability. The information needed for classification is localized to fixed, small patches around lattice sites, meaning that one of CNNs’ primary advantages—spatial invariance—offers little benefit. In fact, enforcing invariance can hinder performance by encouraging the model to search for features that are not physically meaningful in this highly constrained setting.
The challenge is amplified by the difficulty of collecting large, clean training datasets. Neutral atom readout is inherently slow, and acquiring images at scale is time-consuming. The process can also introduce label noise due to state preparation or measurement errors, further reducing the effectiveness of supervised training. As a result, models must learn from limited and imperfect data, unlike in typical vision benchmarks with millions of labeled examples are available.
These limitations become critical at low signal-to-noise ratio (SNR), where the photon count distributions for bright () and dark () states overlap due to noise sources such as photon shot noise and camera readout noise. In this regime, the local intensity patterns that CNNs learn to detect the signal become statistically indistinguishable from random noise fluctuations. Consequently, the network struggles to maintain discriminability between qubit states.
@@ 4. Enabling Readout for Low Photon Counts
In neutral atom systems, qubit readout uses site-resolved fluorescence imaging, where each atom site’s pixel intensity reflects the number of scattered photons collected during a finite illumination period. The detected signal scales linearly with integration time, while dominant noise sources, such as shot noise and camera sensor noise, grow proportionally to the square root of integration time. Therefore, shorter exposures incur a steep drop in SNR, degrading readout fidelity. Longer exposures improve the SNR but risk heating atoms, increasing loss rates and adding latency, thus creating a fundamental trade-off between speed and accuracy.
Figure 5 shows the scaling mismatch enhances the signal relative to background noise more rapidly with increase in readout duration. Long exposures push occupied-site intensities well above the background noise, enabling even simple thresholding to achieve high accuracy. In short exposures, intensities approach the noise floor, causing bright and dark site distributions to overlap and masking spatial features critical for downstream classification tasks.
@@ 4.1. Improving signal-to-noise ratio
In the low-photon regime, effective denoising is essential for recovering and preserving the discriminative features needed to achieve high accuracy in downstream classification. Conventional techniques such as Gaussian smoothing, median filtering, or Wiener filtering can be effective under moderate noise conditions, where they reduce noise by averaging or selectively filtering image components. However, when noise levels approach the signal amplitude these methods often fail. They tend to blur fine spatial details and distort the point-spread patterns that characterize each site’s intensity, thereby removing the critical features required for reliable classification. In such cases, traditional filtering leaves downstream classifiers with insufficient information to accurately distinguish between occupied and unoccupied sites.
Generative approaches offer a more powerful alternative by learning a direct mapping from low-SNR (short exposure) to high-SNR (long exposure) images. Using paired experimental datasets, these models can exploit the repetitive lattice structure of neutral atom arrays, preserve per-site correlations, and selectively enhance meaningful features while suppressing noise, thereby framing the task as an image-to-image translation problem.
Conditional generative adversarial networks (cGANs), such as Pix2Pix, are well-suited for the image-to-image translation problems (Isola et al., 2017). The generator restores missing spatial details, recovers contrast between bright (occupied) and dark (unoccupied) sites, and enhances fine-scale structure, while the discriminator acts as a critic to ensure that the generated outputs are consistently close to the true high-photon-count counterparts as shown in Figure 6(a). These models produce denoised images with the resolution and contrast required for accurate downstream classification in the low-photon regime as shown in Figure 6(b).
Figure 7 shows generative denoising improving low-SNR short-readout images (a) toward the clear site separation of long-exposure images (b), with (c) recovering much of the lost contrast. Figure 7(d) and (e) compare site-averaged intensities with the long-exposure reference. Correctly denoised sites fall in Q1 (bright–bright) and Q3 (dark–dark), whereas Q2 (bright–dark) and Q4 (dark–bright) indicate mismatches. In Figure 7(d), noise scatters many points into Q2 and Q4, but in Figure 7(e) denoising pulls them back into Q1 and Q3, showing closer agreement with ground truth and fewer mismatches. By improving effective SNR without altering the physical readout process, generative denoising directly enhances classification accuracy in low-photon regime, lowering logical error rates and reducing the error-correction overhead in large-scale quantum processors.
@@ 4.2. Deep-dive into generative models for denoising
To meet these requirements, we developed a custom conditional GAN framework, GANDALF (Generative Adversarial Network for Denoising At Low Fluorescence) as shown in Table 2. We employ similar structure as Pix2Pix architecture, but tailor it to learn the mapping from low-SNR to high-SNR images of neutral atom readout. The generator follows an encoder–bottleneck–decoder design with skip connections and residual blocks, combining stability in training with strong feature preservation. Skip connections help mitigate vanishing gradients in deeper networks, allowing early-layer features to directly influence the final reconstruction, which is important for retaining high-frequency spatial detail that might otherwise be lost during downsampling. Residual blocks in the bottleneck further improve feature retention and stability by refining representations without disrupting the preserved information.
@@
where is the low-SNR input, is the high-SNR target, measures how well the generator fools the discriminator, and controls the trade-off between perceptual realism and pixel-level accuracy. In our experiments, provided a good balance between structural preservation and noise suppression. To further improve structural preservation, the objective can be augmented with SSIM (Structural Similarity Index) and PSNR (Peak Signal-to-Noise Ratio) terms, which reward reconstructions that maintain perceptual structure and maximize signal strength relative to noise.
@@ 4.3. Denoising limitations for neutral atom readout
In the low-SNR and low structural variability in readout images, GAN-based denoising can introduce systematic errors despite producing visually cleaner images. Overfitting to recurring noise patterns may distort true signal distributions, and in borderline cases where occupancy depends on subtle intensity differences, the generator may hallucinate sites or suppress faint signals, reducing reconstruction fidelity.
@@ 4.4. Training strategies to mitigate failure modes
Mode collapse can be reduced by preventing the discriminator from becoming overconfident through label smoothing, i.e, assigning real samples a label of 0.9 instead of 1.0 and fake samples 0.1 instead of 0.0. We further reduce collapse and hallucination risk by updating the discriminator less frequently than the generator, giving generator more opportunity to refine outputs without being prematurely constrained.
Dropout in the discriminator encourages focus on robust, generalizable features rather than memorizing noise patterns. To promote smooth convergence and reduce oscillatory dynamics, cosine annealing learning rate schedules are employed. Finally, early stopping based on validation L1 loss captures the best model before it begins overfitting to training-set noise characteristics. Together, these measures allow the model to reliably reconstruct from low-SNR to high-SNR readouts without introducing artificial features, improving classification accuracy in the low-photon regime.
@@ 4.5. Downstream Readout Classification Task
Accurate readout classification is critical for high-fidelity state detection. Traditional thresholding approaches are lightweight but fragile in low-photon regimes. Matched filtering offers interpretability but similarly degrades at low SNR. CNN-based classifiers achieve higher accuracy at the cost of excessive complexity with for atom array (Phuttitarn et al., 2024), making impractical to scale them for larger atom arrays.
GANDALF mitigates this challenge by denoising short-exposure measurements, reducing classification hardness and enabling the use of lightweight models without sacrificing fidelity. The task of neutral atom readout is formulated as binary classification, i.e, given a fixed-size site-centered patch, determine whether the atom is bright-occupied () or dark-unoccupied (). Each of the site-specific models are trained per lattice location to account for spatial variation in fluorescence and noise.
We design and evaluate matched filter with neural network (MF+NN), feedforward neural network (FNN), and 2D-CNN classifiers, to compare their strengths and limitations. Figure 8 illustrates the classification pipeline. The process begins with denoises the readout image with GANDALF, followed by extraction of fixed-size patches centered on each lattice site. CNNs operate directly on these patches, whereas FNN and MF+NN use flattened patches. In case of MF+NN, matched-filter features are additionally incorporated.
@@ 5. Scalable and Fast Readout for QEC
Supporting QEC at scale requires high-fidelity in-circuit measurements across hundreds or thousands of qubits. At such scales, readout speed is a primary bottleneck. While two-qubit gates finish in less than a microsecond, the current readout can take tens of milliseconds. To overcome this gap, two factors are critical: (1) Image acquisition — the duration needed to collect scattered photons. (2) Classification latency — the time required to process the signal and infer qubit states for thousands of qubits. Shorter image acquisition times reduce delay but make classification harder, and often demand more sophisticated models with higher inference costs. In this section, we analyze the speed–accuracy tradeoff and the scalability of GANDALF.
@@ 5.1. Scaling to larger atom arrays
Our approach overcomes this by employing a fully convolutional denoising network, which naturally accepts images of arbitrary size. Convolutional filters operate locally and reuse the same learned kernels across space, allowing the model to generalize seamlessly from small calibration arrays to much larger lattices without modifying architecture or parameters. To evaluate scalability, we test GANDALF on synthetically generated large atom array images. We generate these by stitching single-atom images.
Table 3 reports PSNR, SSIM, and mean L1 loss for denoising as atom arrays scale from to . The results show only slight degradation with increasing grid size, i.e, PSNR drops by less than 0.5dB, SSIM decreases by 0.02, and mean L1 loss remains nearly unchanged. With fixed Model complexity, the computational cost of GANDALF grows marginally with image size, keeping per-site inference time nearly constant.
| Metric | Grid Size () | |||
@@ 5.2. Denoising Latency and Practical Scalability
We evaluate the GANDALF generator (4.7M parameters) across grid sizes, batch sizes, and parallel executions to assess its suitability for real-time QEC pipelines. Although input size affects inference time, the fully convolutional design ensures sublinear growth on a single GPU, with latency remaining within tens of ms even for arrays as shown in Table 3. These trends indicate that denoising is practical at scale, with GPUs or other accelerators.
@@
Throughput can be further improved using batching and parallelism. With batching, the GPU processes multiple images simultaneously, reducing per-image latency from ms to ms at a batch size of 128 (Table 4). Parallel execution provides another dimension of scaling: as shown in Table 5, throughput per model remains nearly constant with latencies of 80–90 s, while memory usage increases roughly linearly with the number of models. In practice, moderate batching combined with 8–32 concurrent models keeps denoising well within QEC timing budgets while sustaining high throughput. Because throughput aggregates and scales nearly linearly with the number of models, practical large-scale readout is achievable even on commodity GPU.

### v2

@@ 3.1. Experimental setup and dataset
We utilize the dataset generated by the neutral atom system consisting of grid of cesium atoms, each held in place by optical tweezers as mentioned in prior work (phuttitarn2023enhanced). To read out the state of each atom, we use a laser beam that causes atoms in the “bright” state to fluorescence, while the “dark” state atoms emit few to no photons. The emitted photons are then captured by a sensitive camera through two optical paths: (1) a bright, high-quality path used for generating accurate labels, and (2) a dimmer, noisier path used for generating readout images for testing machine learning models.
The experimental setup is shown in Figure 3 is used to collected the readout images of grid of Cesium (Cs) atom arrays resulting in single-site images, with exposure durations ranging from to for primary path and from to for secondary path with noisy images. The atom presence labels are generated using primary path using thresholding methods and the noisy images are used for testing machine learning models.
3.2. Low photon counts limits fast readout
Each readout image obtained by the EMCCD camera consists of fluorescence from atoms arranged on a fixed lattice, where each site is either bright () or dark (). Fast neutral-atom readout is fundamentally limited by the number of fluorescence photons collected during a measurement. Photon arrivals follow Poisson statistics,666The number of detected photons in a given exposure has mean and variance , yielding a fundamental shot-noise level of . so, while the signal grows linearly with exposure time , the signal-to-noise ratio (SNR) improves only as . As shown in Figure 4, short exposures (1.5–10 ms) collect far fewer photons and produce signals that barely exceed the background noise floor, whereas longer exposures (15–100 ms) accumulate enough photons to clearly separate occupied and unoccupied sites. This scaling of signal-to-noise ratio makes aggressive reductions in exposure time rendering fast readout inherently challenging.
3.3. Downstream classification challenges
The reduction in photon counts at short exposures also significantly complicates downstream state classification. As shown in Figure 5(b), long-exposure images yield well-separated intensity distributions for bright and dark sites, allowing simple classifiers, including thresholding with Gaussian or square masks and Gaussian mixture models (GMMs), to perform reliably. At short exposures, however, the distributions for the two states overlap substantially due to insufficient photon counts as shown in Figure 5(a), eliminating the contrast required for these lightweight methods to discriminate states accurately. This loss of separability necessitates more expressive downstream machine learning classifiers capable of extracting discriminative features from weak, noise-dominated short exposure readout images.
Prior work (phuttitarn2023enhanced) developed two CNN-based classifiers that substantially improve readout accuracy relative to thresholding approaches. These CNNs are trained and tested on the low-exposure noisy images with the labels generated from the high-quality optical path. CNN-site (355k parameters) operates on individual lattice sites and reduces classification errors by up to 56% in low-crosstalk settings, while CNN-array (74M parameters) incorporates neighboring-site information and improves error rates by up to 43% and crosstalk misclassification by 78% compared to Gaussian thresholding.
Although these models meaningfully extend the feasible operating range to moderately shorter exposures, their accuracy remains limited when photon counts are very low, as they must classify directly from heavily degraded images with minimal signal contrast. Further gains would require even larger and more complex CNNs, an approach that is both ineffective—because the underlying signal is too weak—and impractical due to the growing model size, resulting in long inference times for larger arrays. This necessitates a more scalable strategy that alleviates the burden on classifiers.
@@ 3.4. Why are standard CNN models not enough?
CNNs are the de-facto models for image classification. In natural-image domains with rich semantic content and large training datasets, they perform remarkably well, even in the presence of moderate image noise, label noise, and variations in scale or rotation, as they learn robust, hierarchical777CNNs operate by first detecting local, low-level features such as edges, corners, and textures in early layers. Deeper layers combine these primitives into higher-level features that capture complex shapes and semantic structure. This feature hierarchy allows CNNs to distinguish objects even under degraded imaging conditions, provided there is enough semantic variation in the data. representations of visual features.
The readout images lack semantic structure and variability and differ fundamentally from natural-image datasets. The information needed for classification is localized to fixed, small patches around lattice sites, meaning that one of CNNs’ primary advantages—spatial invariance—offers little benefit. In fact, enforcing invariance can hinder performance by encouraging the model to search for features that are not physically meaningful in this highly constrained setting.
The challenge is amplified by the difficulty of collecting large, clean training datasets. Neutral atom readout is inherently slow, and acquiring images at scale is time-consuming. The process can also introduce label noise due to state preparation or measurement errors, further reducing the effectiveness of supervised training. As a result, models must learn from limited and imperfect data, unlike in typical vision benchmarks for which millions of labeled examples are available. These limitations become critical in a low signal-to-noise ratio (SNR) regime888The photon count distributions for bright () and dark () states overlap due to noise sources such as photon shot noise and camera readout noise., where the local intensity patterns that CNNs learn to detect the signal become statistically indistinguishable from random noise fluctuations. Consequently, the network struggles to maintain discrimination accuracy between qubit states.
@@ 4. Enabling readout for low photon counts
Neutral atom readout uses site-resolved fluorescence imaging: each atom sits at a known location in the image, and the brightness of the corresponding pixel tells us how many photons we collected during a short illumination period. If we make this readout shorter, we collect fewer photons. The useful signal then shrinks in direct proportion to the readout time, while the random noise shrinks more slowly (roughly with the square root of time). This makes it progressively harder to tell bright (occupied) sites from dark (empty) ones as we speed up the measurement, creating a fundamental trade-off between readout speed and accuracy.
Standard image-processing methods—such as Gaussian smoothing, median filtering, or Wiener filtering—work reasonably well when the images are only moderately noisy. They can reduce fluctuations while keeping the overall pattern intact. In the low-photon regime, however, the noise becomes comparable to the signal itself. These filters then tend to blur sharp features and distort the characteristic light pattern at each site, removing exactly the details needed to decide whether a site is occupied or empty. In this regime, we need denoising methods that can clean up the image while preserving these subtle, discriminative features so that downstream classifiers can still achieve high accuracy.
@@ 4.1. Improving signal-to-noise ratio
Generative approaches offer a more powerful alternative to conventional techniques by learning a direct mapping from low-SNR (short exposure) to high-SNR (long exposure) images. Using paired experimental datasets, these models can exploit the repetitive lattice structure of neutral atom arrays, preserve per-site correlations, and selectively enhance meaningful features while suppressing noise, framing the task as image-to-image translation problem.
Conditional generative adversarial networks (cGANs), such as Pix2Pix, are well-suited for image-to-image translation problems (pix2pix2017). The generator restores missing spatial details, recovers contrast between bright (occupied) and dark (unoccupied) sites, and enhances fine-scale structure, while the discriminator acts as a critic to ensure that the generated outputs, the denoised images, are consistently close to the true high-photon-count counterparts as shown in Figure 6(a). These models produce denoised images with the resolution and contrast required for accurate downstream classification in the low-photon regime as shown in Figure 6(b).
Generative denoising improves low-SNR short-readout images by generating denoised images that are closer to the high-SNR long-exposure images by recovering much of the lost contrast. The noisy, low-SNR short exposure readout image in Figure 7(a) is denoised to Figure 7(c) that resembles the high-SNR long-exposure image in Figure 7(b).
To evaluate this quantitatively, Figure 7(d - e) compares site-averaged intensities with the long-exposure reference for a particular site for a group of site-specific images. Ideally, we want to observe a diagonal line with minimal deviations. Correctly denoised sites fall in Q1 (bright–bright) and Q3 (dark–dark), whereas Q2 (bright–dark) and Q4 (dark–bright) indicate mismatches. In Figure 7(d), noise scatters many points into Q2 and Q4, but in Figure 7(e) denoising pulls them back into Q1 and Q3, showing closer agreement with ground truth and fewer mismatches as expected. By improving effective SNR without altering the physical readout process, generative denoising directly enhances classification accuracy in low-photon regime, lowering logical error rates and reducing the error-correction overhead in large-scale quantum processors.
@@ 4.2. Deep-dive into generative models for denoising
To meet these requirements, we developed a custom conditional GAN framework, GANDALF (Generative Adversarial Network for Denoising At Low Fluorescence) as shown in Table 2. We employ a similar structure as the Pix2Pix architecture, but tailor it to learn the mapping from low-SNR to high-SNR images of neutral atom readout. The generator follows an encoder–bottleneck–decoder design with skip connections and residual blocks, combining stability in training with strong feature preservation. Skip connections help mitigate vanishing gradients in deeper networks, allowing early-layer features to directly influence the final reconstruction, which is important for retaining high-frequency spatial detail that might otherwise be lost during downsampling. Residual blocks in the bottleneck further improve feature retention and stability by refining representations without disrupting the preserved information.
@@
where is the low-SNR input, is the high-SNR target, measures the ability of generator to fool the discriminator, as mentioned in (pix2pix2017), and controls the trade-off between perceptual realism and pixel-level accuracy. In our experiments, provided a good balance between structural preservation and noise suppression. To further improve structural preservation, the objective can be augmented with the Structural Similarity Index Measure (SSIM) and Peak Signal-to-Noise Ratio (PSNR) terms, which reward reconstructions that maintain perceptual structure and maximize SNR.
@@ 4.3. Denoising limitations for neutral atom readout
In the low-SNR regime and with low structural variability in readout images, GAN-based denoising can introduce systematic errors despite producing visually cleaner images. Overfitting to recurring noise patterns may distort true signal distributions, and in borderline cases where occupancy depends on subtle intensity differences, the generator may hallucinate sites or suppress faint signals, reducing reconstruction fidelity.
@@ 4.4. Training strategies to mitigate failure modes
Mode collapse can be reduced by preventing the discriminator from becoming overconfident through label smoothing, i.e, assigning real samples a label of 0.9 instead of 1.0 and fake samples 0.1 instead of 0.0. We further reduce collapse and hallucination risk by updating the discriminator less frequently than the generator, giving the generator more opportunities to refine outputs without being prematurely constrained.
Dropout in the discriminator encourages focus on robust, generalizable features rather than memorizing noise patterns. To promote smooth convergence and reduce oscillatory dynamics, cosine annealing learning rate schedules are employed. Finally, early stopping based on validation L1 loss captures the best model before it begins overfitting to training-set noise characteristics. Together, these measures allow the model to reliably reconstruct from low-SNR to high-SNR readouts without introducing artificial features, improving classification accuracy in the low-photon-count regime.
@@ 4.5. Downstream Readout Classification Task
Accurate readout classification is critical for high-fidelity state detection. Traditional thresholding approaches are lightweight but fragile in low-photon regimes. Matched filtering offers interpretability but similarly degrades at low SNR. Convolutional Neural Networks (CNN)-based classifiers achieve higher accuracy at the cost of excessive complexity with million parameters for a atom array (phuttitarn2023enhanced), making it impractical to scale them for larger atom arrays.
GANDALF mitigates this challenge by denoising short-exposure measurements, reducing classification hardness and enabling the use of lightweight models without sacrificing fidelity. The task of neutral atom readout is formulated as binary classification. Given a fixed-size site-centered patch, determine whether the atom is bright-occupied () or dark-unoccupied (). Each site-specific model is trained per lattice location to account for spatial variations in signal and noise.
We design and evaluate matched filter with neural network (MF+NN), feedforward neural network (FNN), and 2D-CNN classifiers, to compare their strengths and limitations. Figure 8 illustrates the classification pipeline. The process begins with denoising the readout image with GANDALF, followed by extraction of fixed-size patches centered on each lattice site. CNNs operate directly on these patches, whereas FNN and MF+NN use flattened patches. In the case of MF+NN, matched-filter features are additionally incorporated.
@@ 5. Scalable and Fast Readout for QEC
Supporting QEC at scale requires high-fidelity in-circuit measurements across hundreds or thousands or even more qubits. At such scales, readout speed is a primary bottleneck. While two-qubit gates finish in less than a microsecond, the current readout can take tens of milliseconds. To overcome this gap, two factors are critical: (1) Image acquisition — the duration needed to collect scattered photons. (2) Classification latency — the time required to process the signal and infer qubit states for thousands of qubits. Shorter image acquisition times reduce delay but make classification harder, and often demand more sophisticated models with higher inference costs. In this section, we analyze the speed–accuracy tradeoff and the scalability of GANDALF.
@@ 5.1. Scaling to larger atom arrays
Our approach overcomes this by employing a fully convolutional denoising network, which naturally accepts images of arbitrary size. Convolutional filters operate locally and reuse the same learned kernels across space, allowing the model to generalize seamlessly from small calibration arrays to much larger lattices without modifying architecture or parameters. To evaluate scalability, we test GANDALF on synthetically generated large atom arrays by stitching single-atom images.
Table 3 reports Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and mean L1 loss for denoising as atom arrays scale from to . The results show only slight degradation with increasing grid size, i.e, PSNR drops by less than 0.5dB, SSIM decreases by 0.02, and mean L1 loss remains nearly unchanged. With fixed model complexity, the computational cost of GANDALF grows marginally with image size, keeping per-site inference time nearly constant.
| Metric | Atom array grid size () | |||
@@ 5.2. Denoising Latency and Practical Scalability
We evaluate the GANDALF generator (4.7 million parameters) across grid sizes, batch sizes, and parallel executions to assess its suitability for real-time QEC pipelines. Although input size affects inference time, the fully convolutional design ensures sublinear growth on a single GPU, with latency remaining within tens of ms even for atom arrays as shown in Table 3. These trends indicate that denoising is practical at scale, with GPUs or other accelerators.
@@
Throughput can be further improved using batching and parallelism. With batching, the GPU processes multiple images simultaneously, reducing per-image latency from ms to ms at a batch size of 128, as shown in Table 4, for an atom array of , similar to the dataset used (phuttitarn2023enhanced).
Parallel execution provides another dimension of scaling, as shown in Table 5, throughput per model remains nearly constant with latencies of 80–90 s, while memory usage increases roughly linearly with the number of models. In practice, moderate batching combined with 8–32 concurrent models keeps denoising well within QEC timing budgets while sustaining high throughput. We envision further reductions in denoising latency by using more powerful GPUs, as our GPU setup has modest computational capabilities and 5x lower memory bandwidth than NVIDIA® H100 GPUs.